OpenStudy (anonymous):

m =y2-1y/x2-x1 for for m = 1 and (x1, y1) = (0, 2) and (x2, y2) = (x, 0). help me :(

OpenStudy (anonymous):

What is the question exactly ?

OpenStudy (anonymous):

to evalute the problem

OpenStudy (anonymous):

m =y2-1y/x2-x1 putting calues we get 1 = 0-2/x-0 1 = -2/x x = -2 is the solution